# Stable Matchings with Choice Correspondences Under Acyclicity

### 24 Mar 2026 v1

@@ Abstract
We study the existence of stable matchings when agents have choice correspondences instead of preference relations. We extend the framework of Chambers and Yenmez (2017) by weakening the Path Independence assumption. For many-to-many markets, we show that stable matchings exist when choice correspondences satisfy Substitutability and a new General Acyclicity condition. We provide a constructive proof using a Grow or Discard Algorithm that iteratively expands or eliminates contracts until a strongly maximal Individually Rational set is reached. We provide an algorithm to obtain stable matchings in which rejected contracts are not permanently discarded, distinguishing our approach significantly from standard DAA-type algorithms. For one-to-one markets, we show that Path Independence alone does not guarantee stability. We introduce a replacement-based notion of stability and provide an algorithm that constructs stable matchings when choice correspondences satisfy Binary Acyclicity.
@@ 1 Introduction
We show that our matching algorithm, the Dynamic Grow or Discard Matching Algorithm (GDMA)) can be implemented where agents on one side of the market propose contracts for the side to accept or reject. Agents on the other side again use the ‘grow or discard’ algorithm for individual choices to decide whether to ‘grow’ by expanding the set or to ‘discard’ some old contracts and move to a different set of contracts. Unlike standard deferred acceptance algorithms, the Dynamic GDMA allows firms to re-propose previously rejected contracts. Proving the stability of the obtained matching, therefore, requires new arguments than the ones used in the literature.
@@
We then provide results for the one-to-one matching problem, which is not studied in CY. We show that Path Independence alone is not sufficient to guarantee stability in the one-to-one setting. To address this, we introduce a notion of stability based on the replacement of current matches when choices are acyclic over binary sets. Our stability notion captures the idea that a matching remains stable not because agents possess a complete ranking, but because no agent can find an alternative partner who would simultaneously choose to replace their current match. We provide a review of the literature surrounding this work to place our contribution.
@@ 4 The many-to-many matching problem
    No CY-blocking: A matching $\nu$ has No CY-blocking if there does not exist a non-empty set of contracts $X\subseteq\mathbf{X}$, with $X\cap\nu=\emptyset$, such that:
@@
A matching $\nu$ is IR if every agent wants to keep all the contracts provided to her in the matching. A matching $\nu$ has No CY-blocking if there is no non-empty set of new contracts $X$ (disjoint from $\nu$) such that, when each firm and worker is offered its current contracts together with its contracts in $X$, every agent would choose to accept all of its contracts in $X$. In other words, there is no collection of contracts that all involved firms and workers would jointly accept in addition to their current match.
@@
[…]
Note that for $k=2$, GA holds vacuously. A violation of GA for $k=2$ would require that $C(S_{1}\cup S_{2})=S_{2}\cup D_{1}$ where $D_{1}\subsetneq S_{1}$, and $C((S_{2}\cup D_{1})\cup S_{1})=S_{1}\cup D_{2}$ where $D_{2}\subsetneq S_{2}\cup D_{1}$. However, this implies that $(S_{1}\cup S_{2})\subseteq C(S_{1}\cup S_{2})$. This along with the fact that $D_{1}\subsetneq S_{1}$ is a contradiction. Therefore, GA prevents cycles of length three and above from occurring in a ‘grow’ or ‘discard’ manner. At each step starting from the first: where we start with the set $S_{1}$ move to another set by adding $S_{2}$ and discarding some elements from $S_{1}$, then moving to $S_{3}$ and discarding some elements from $S_{2}\cup D_{1}$, and so on, ending with the last stage where we again obtain the whole set $S_{1}$ which was (partially) ‘discarded’ before reaching the last stage. We now show that PI implies GA.
@@
Let $X=\{a,b,c,d\}$. Let the choices be as shown Table 4. We note that the choice correspondence does not have a ‘stable’ IR set even though it satisfies SUB. For example, even though $\{a,b,d\}$ is IR, $C(\{a,b,d\}\cup\{c\})=C(X)=\{a,d\}$. Therefore, property (ii) mentioned above is not satisfied for any IR set. Similarly, singleton sets are also not stable since $C(a,b)=\{a,b\}$ (implies neither $\{a\}$ nor $\{b\}$ are stable, $C(c,d)=d$ ($\{c\}$ is not stable), $C(a,d)=\{a,d\}$ ($\{d\}$ is not stable).
@@
Theorem 1 provides weaker sufficient conditions on choice correspondences for the existence of CY-stable matchings.
The proof is provided in the Appendix. We provide a sketch of the proof. First, we show that there must exist a strongly maximal IR set as defined earlier using a Grow or Discard Algorithm (GDA) for individual choice correspondences. We then use the Dynamic Grow or Discard Matching Algorithm (GDMA) to arrive at a stable matching through proposals and rejections. This algorithm is significantly different from DAA since rejected contracts are not discarded permanently by the proposing side. We prove the termination of the algorithm using finiteness of the set of all contracts and GA.
We present the GDA (Algorithm 1) in the table below which always produces a strongly maximal IR set. [^18] See the proof for a formal definition. It starts with $C(\mathcal{X})$ and proceeds to add a disjoint set to it if indicated by the choice correspondence.
@@ 6 Conclusion
This matching is individually rational by construction, as each firm only proposes contracts to a worker which is individually rational for the firm. Consider a blocking contract $x$ which blocks the matching in the sense of stability, in the firm proposing DAA. Denote by $(f,w)$ the firm and worker pair associated with $x$. It must be that $C_{f}(\{\nu_{f},x\})=x$ and $C_{w}(\{\nu_{w},x\})=x$ where $\nu_{f}$ is $\nu_{w}$’s are the contracts matched to the worker and the firm in the firm proposing DAA. But then $f$ must have proposed the contract $x$ first since it is individually rational as it is preferred to $\nu_{f}$ and been rejected in favor of a preferred match $x^{\prime}$ which is rejected by $w$ in favor of $x^{\prime}$, which in turn gets rejected in favor of $x^{\prime\prime}$ and this sequence eventually stops when the worked accepts $\nu_{w}$ and does not reject it. By BA, we then have $x\in c_{w}(\{x^{\prime},\nu_{w}\})$, which is a contradiction. Thus, such a matching will be stable. ∎
Proof of Theorem 1
Proof.
[…]
Definition 10 (Maximality).
For any set $X\in\mathcal{X}_{f}$,
Lemma 4.
There exists at least one strongly maximal set which is IR.
Proof.
Lemma 5.
Proof.
$C_{f}(X_{1})=X_{1}=X_{t},$
$C_{f}(X_{1}\cup X_{2})=X_{2}\cup D_{1}=X_{t+1},D_{1}\subseteq X_{1},$
$C_{f}((X_{2}\cup D_{1})\cup X_{3})=X_{3}\cup D_{2},D_{2}\subseteq X_{2}\cup D_{1},$
$\vdots$
$C_{f}(X_{1}^{*})=X_{1}^{*}=X_{t}^{*},$
$\vdots$
Similar properties can be shown for $C_{w}(.)$ for every $w\in\mathbf{W}$. The firm proposing Dynamic Grow or Discard Matching Algorithm (GDMA) was presented on page 18. We show that the matching algorithm will produce a stable matching. Let $\nu$ be the final matching produced when the GDMA terminates. To prove that $\nu$ is a stable matching, we must show that it is individually rational (IR) and that there exists no blocking coalition.
Step 1: Termination
Step 2: Absence of a Blocking Coalition
  1. 1.
    $Z_{f}\subseteq C_{f}(\nu_{f}\cup Z_{f})$ for all firms $f$ involved in $Z$, and
  2. 2.
    $Z_{w}\subseteq C_{w}(\nu_{w}\cup Z_{w})$ for all workers $w$ involved in $Z$.
Step 3: The Terminal State
Therefore, the assumption that a blocking coalition $Z$ exists must be false. Since $\nu$ is individually rational and admits no blocking coalitions, the matching $\nu$ produced by the GDMA is stable. ∎

### 30 Mar 2026 v3

@@ Abstract
We study the existence of stable matchings when agents have choice correspondences instead of preference relations. We extend the framework of Chambers and Yenmez (2017) by weakening the Path Independence assumption. For many-to-many markets, we show that stable matchings exist when choice correspondences satisfy Substitutability and a new General Acyclicity condition. We provide a constructive proof using a Grow or Discard Algorithm that iteratively expands or eliminates contracts until a strongly maximal Individually Rational set is reached. We provide an algorithm to obtain stable matchings in which rejected contracts are not permanently discarded, distinguishing our approach significantly from standard DAA-type algorithms. For one-to-one markets, we introduce a replacement-based notion of stability and provide an algorithm that constructs stable matchings when choice correspondences satisfy Binary Acyclicity, a property weaker than Path Independence.
@@ 1 Introduction
We show that our matching algorithm, the Dynamic Grow or Discard Matching Algorithm (GDMA) can be implemented where agents on one side of the market propose contracts for the side to accept or reject. Agents on the other side again use the ‘grow or discard’ algorithm for individual choices to decide whether to ‘grow’ by expanding the set or to ‘discard’ some old contracts and move to a different set of contracts. Unlike standard Deferred Acceptance Algorithms, the Dynamic GDMA allows firms to re-propose previously rejected contracts. Proving the stability of the obtained matching, therefore, requires new arguments than the ones used in the literature.
@@
We then provide results for the one-to-one matching problem, which is not studied in CY. We introduce a different notion of stability based on the replacement of current matches when choices are acyclic over binary sets. Our stability notion captures the idea that a matching remains stable not because agents possess a complete ranking, but because no agent can find an alternative partner who would simultaneously choose to replace their current match. We provide a review of the literature surrounding this work to place our contribution.
@@ 4 The many-to-many matching problem
    No blocking: A matching $\nu$ has No blocking if there does not exist a non-empty set of contracts $X\subseteq\mathbf{X}$, with $X\cap\nu=\emptyset$, such that:
@@
A matching $\nu$ is IR if every agent wants to keep all the contracts provided to her in the matching. A matching $\nu$ has No blocking if there is no non-empty set of new contracts $X$ (disjoint from $\nu$) such that, when each firm and worker is offered its current contracts together with its contracts in $X$, every agent would choose to accept all of its contracts in $X$. In other words, there is no collection of contracts that all involved firms and workers would jointly accept in addition to their current match.
@@
We can use the notion of revealed preference through choice correspondence as follows: $A$ is revealed preferred to $B$ if $A\subseteq C(A\cup B)$. GA requires that revealed preferred cycles of the following kind do not occur: $S_{2}$ is revealed preferred to $S_{1}$, $S_{3}$ is revealed preferred to the choice made in the previous stage, (i.e. $S_{2}\cup D_{1}$) and so on till $S_{k-1}$ is revealed preferred to $(S_{k-2}\cup D_{k-3})$ and finally $S_{1}$ (through $S_{k}$) is revealed preferred to $S_{k-1}\cup D_{k-1}$. [^15] Note that no relationship is assumed between $S_{1}$ and $S_{k}$. In other words, choice over pairwise disjoint sets should not ‘cycle’ in a way where the initial set is discarded and then again selected (as a subset possibly) at the end of the cycle.
Note that for $k=2$, GA holds vacuously. A violation of GA for $k=2$ would require that $C(S_{1}\cup S_{2})=S_{2}\cup D_{1}$ where $D_{1}\subsetneq S_{1}$, and $C((S_{2}\cup D_{1})\cup S_{1})=S_{1}\cup D_{2}$ where $D_{2}\subsetneq S_{2}\cup D_{1}$. However, this implies that $(S_{1}\cup S_{2})\subseteq C(S_{1}\cup S_{2})$. This along with the fact that $D_{1}\subsetneq S_{1}$ is a contradiction. Therefore, GA prevents cycles of length three and above from occurring in a ‘Grow’ or ‘Discard’ manner. At each step starting from the first: where we start with the set $S_{1}$ move to another set by adding $S_{2}$ and discarding some elements from $S_{1}$, then moving to $S_{3}$ and discarding some elements from $S_{2}\cup D_{1}$, and so on, ending with the last stage where we again obtain the whole set $S_{1}$ which was (partially) discarded before reaching the last stage. We now show that PI implies GA.
@@
Let $X=\{a,b,c,d\}$. Let the choices be as shown Table 4. We note that the choice correspondence does not have a stable IR set even though it satisfies SUB. For example, even though $\{a,b,d\}$ is IR, $C(\{a,b,d\}\cup\{c\})=C(X)=\{a,d\}$. Therefore, property (ii) mentioned above is not satisfied for any IR set. Similarly, singleton sets are also not stable since $C(a,b)=\{a,b\}$ (implies neither $\{a\}$ nor $\{b\}$ are stable, $C(c,d)=d$ ($\{c\}$ is not stable), $C(a,d)=\{a,d\}$ ($\{d\}$ is not stable).
@@
Theorem 1 establishes that the classical requirement of Path Independence can be safely relaxed, providing strictly weaker sufficient conditions than those in the existing literature for the existence of CY-stable matchings.
The formal proof is relegated to the Appendix; here, we outline the underlying intuition. The proof proceeds in two main steps. First, we establish the existence of a strongly maximal individually rational (IR) set for individual choice correspondences via a novel Grow or Discard Algorithm (GDA). Second, we deploy the Dynamic Grow or Discard Matching Algorithm (GDMA) to simulate the market’s proposal and rejection dynamics until a stable matching is reached. Crucially, the GDMA departs fundamentally from the classical Deferred Acceptance Algorithm (DAA): because agents are not bound by the Irrelevance of Rejected Contracts, proposing agents do not permanently discard rejected contracts. Instead, the algorithm successfully navigates this path-dependent choice behavior to converge on stability.
We now present the GDA (Algorithm 1) in the table below which always produces a strongly maximal IR set for an agent’s choice correspondence. [^18] See the proof for a formal definition. It starts with $C(\mathcal{X})$ and proceeds to add a disjoint set to it if indicated by the choice correspondence.
@@ 6 Conclusion
[…]
Definition 10 (Maximality).
For any set $X\in\mathcal{X}_{f}$,
Lemma 3.
There exists at least one strongly maximal set which is IR.
Proof.
Lemma 4.
Proof.
$C_{f}(X_{1})=X_{1}=X_{t},$
$C_{f}(X_{1}\cup X_{2})=X_{2}\cup D_{1}=X_{t+1},D_{1}\subseteq X_{1},$
$C_{f}((X_{2}\cup D_{1})\cup X_{3})=X_{3}\cup D_{2},D_{2}\subseteq X_{2}\cup D_{1},$
$\vdots$
$C_{f}(X_{1}^{*})=X_{1}^{*}=X_{t}^{*},$
$\vdots$
Similar properties can be shown for $C_{w}(.)$ for every $w\in\mathbf{W}$. The firm proposing Dynamic Grow or Discard Matching Algorithm (GDMA) was presented on page 18. We show that the matching algorithm will produce a CY-stable matching. Let $\nu$ be the final matching produced when the GDMA terminates. To prove that $\nu$ is a CY-stable matching, we must show that it is Individually Rational (IR) and that there exists no blocking coalition.
Step 1: Termination
Step 2: Absence of a Blocking Coalition
  1. 1.
    $Z_{f}\subseteq C_{f}(\nu_{f}\cup Z_{f})$ for all firms $f$ involved in $Z$, and
  2. 2.
    $Z_{w}\subseteq C_{w}(\nu_{w}\cup Z_{w})$ for all workers $w$ involved in $Z$.
Therefore, the assumption that a blocking coalition $Z$ exists must be false. Since $\nu$ is Individually Rational and admits no blocking coalitions, the matching $\nu$ produced by the GDMA is CY-stable. ∎
Proof of Theorem 2
Proof.
@@
This matching is Individually Rational by construction, as each firm only proposes contracts to a worker which is Individually Rational for the firm. Consider a blocking contract $x$ which blocks the matching in the sense of stability, in the firm proposing DAA. Denote by $(f,w)$ the firm and worker pair associated with $x$. It must be that $C_{f}(\{\nu_{f},x\})=x$ and $C_{w}(\{\nu_{w},x\})=x$ where $\nu_{f}$ is $\nu_{w}$’s are the contracts matched to the worker and the firm in the firm proposing DAA. But then $f$ must have proposed the contract $x$ first since it is Individually Rational as it is preferred to $\nu_{f}$ and been rejected in favor of a preferred match $x^{\prime}$ which is rejected by $w$ in favor of $x^{\prime}$, which in turn gets rejected in favor of $x^{\prime\prime}$ and this sequence eventually stops when the worked accepts $\nu_{w}$ and does not reject it. By BA, we then have $x\in c_{w}(\{x^{\prime},\nu_{w}\})$, which is a contradiction. Thus, such a matching will be R-stable. ∎